OpenStudy (anonymous):

with approaximately what velocity do you enter the water if you dive from a 10 metre platform

OpenStudy (anonymous):

\[Vf^2=Vi^2+2ad, Vi=0 \because free fall, a=9.8m/s^2 gravity, d=10m\]